### 15 Apr 2026 v1

# Tinniam V Ganesh

Abstract
This paper develops a unified Markov Decision Process (MDP) framework for optimising two recurring in-match decisions in T20 cricket viz. batting order selection and bowling plan assignment, directly in terms of win and defend probability rather than expected runs. A three-phase player profile engine (Powerplay, Middle, Death) with James–Stein shrinkage is estimated from 1,161 IPL ball-by-ball records (2008–2025). Win/defend probabilities are evaluated by vectorised Monte Carlo simulation over N=50, [^0] N=50{,}000 innings trajectories. Batting orders are searched by exhaustive enumeration. Bowling plans are computed by simulated annealing over the remaining quota with the constraint that the same bowler cannot bowl consecutive overs. Applied to two 2026 IPL matches, the optimal batting order improves Mumbai Indians’ win probability by 4.1 pp (52.4% → 56.5%), and the optimal Gujarat Titans bowling plan improves defend probability by 5.2 pp (39.1% → 44.3%). In both cases the source of suboptimality is phase-agnostic deployment of players.
Keywords: T20 cricket; batting order optimisation; bowling plan optimisation; Markov Decision Process; Monte Carlo simulation; simulated annealing; James–Stein shrinkage; sports analytics; Indian Premier League
1. Introduction
The Twenty20 (T20) format has compressed professional cricket into 120 deliveries per side, yet within that frame a single over can determine the outcome. In the Indian Premier League (IPL), where franchise valuations exceed one billion USD and playoff positions hinge on net run rate, the quality of two recurring tactical decisions is decisive: which batsman to promote at each wicket fall, and which bowler to assign to each remaining over.
Both decisions are sequential and state-dependent. The batting captain must balance aggression against preservation given the current match state (runs remaining, balls remaining, wickets in hand) and the batsmen still available. The bowling captain must assign bowlers subject to the T20 quota (maximum four overs per bowler) and the consecutive-over rule, while accounting for phase-specific economy and wicket-taking ability. Both decisions are therefore naturally modelled as Markov Decision Processes (Bellman, 1957; Puterman, 1994), with the objective of maximising a probability — that runs scored cross a target, or that runs conceded stay below it — rather than an expectation. Prior linear programming approaches (Ganesh, 2017) optimise expected runs, but maximising expected runs does not maximise win probability — especially when the match is close and a team needs a specific number of runs rather than simply as many as possible.
This paper makes the following contributions:
  1. 1.
    A unified MDP framework that models batting order and bowling plan optimisation on a common state space with symmetric Bellman equations, making explicit the duality between the batting win probability and the bowling defend probability.
  2. 2.
    A three-phase player profile engine using Laplace smoothing and James–Stein shrinkage that estimates per-ball outcome distributions for batsmen and bowlers in Powerplay (overs 1–6), Middle (overs 7–15), and Death (overs 16–20) phases separately, from 1,161 historical IPL matches.
  3. 3.
    Simulation-based search, namely exhaustive enumeration for batting order and simulated annealing for bowling plans, that directly optimises win/defend probability over the stochastic innings model, validated at N=50, [^0] N=50{,}000 simulations per configuration (Monte Carlo SE $\approx 0.22\%$).
  4. 4.
    Two detailed case studies from the 2026 IPL season, providing quantitative audits of actual in-match decisions: the batting and bowling decisions in KKR vs. MI (29 March 2026) and the bowling decisions in GT vs. PBKS (31 March 2026). The two cases span different match contexts: one where the bowling gap is large (+18.1 pp) and one where it is modest (+5.2 pp), enabling a comparative analysis of the conditions under which sub-optimal decisions have the greatest impact.
The paper is structured as follows. Section 2 reviews related work. Section 3 presents the MDP framework. Section 4 describes the player profile engine. Section 5 presents the Monte Carlo simulation. Section 6 covers the search algorithms. Sections 7 and 8 present the two case studies and their results. Section 9 provides a comparative analysis and discussion of limitations. Section 10 concludes.
2. Related Work
The Duckworth–Lewis method (Duckworth and Lewis, 1998) implicitly defines a resource table over (runs needed, balls remaining, wickets in hand), the same state space used in this paper, for resetting targets in interrupted matches. It remains the de facto standard in international cricket and establishes that match state can be meaningfully quantified.
The author’s earlier work (Ganesh, 2017) formulated batting lineup and bowling selection as linear programming problems, assigning ball-level encounters $o_{ij}$ between player $i$ and bowler $j$ to maximise aggregate strike rate (batting) or minimise aggregate economy rate (bowling). That formulation correctly identifies that the allocation of playing time across player matchups matters, but it optimises expected runs rather than win probability and cannot encode the sequential, constraint-rich structure of T20 over assignment. The present paper replaces the LP objective with a stochastic MDP framework whose objective directly matches the match outcome: a win probability threshold rather than an expectation.
The MDP framework and Bellman equations are due to Bellman (1957) and Puterman (1994). The James–Stein shrinkage estimator applied to sparse player profiles follows James and Stein (1961). The simulated annealing search for bowling plans draws on Kirkpatrick, Gelatt, and Vecchi (1983).
3. The MDP Framework
3.1 State Space and Match Dynamics
Let the match state at any ball be the triple:
$\mathcal{S}=(r,\,b,\,w)$
where $r\in\{1,\ldots,R_{\max}\}$ is the runs remaining for the batting side to win (or, equivalently, for the bowling side to defend), $b\in\{0,1,\ldots,B\}$ is the number of legal balls remaining in the innings ($B=120$ in T20), and $w\in\{0,1,\ldots,10\}$ is the number of wickets in hand (batting perspective) or wickets taken (bowling perspective). The Markov property holds: the probability distribution over future match trajectories is fully determined by $\mathcal{S}$, independent of the history of play.
On each legal delivery, one of seven outcomes $o\in\Omega=\{W,0,1,2,3,4,6\}$ is realised, where $W$ denotes the dismissal of the on-strike batsman. The runs-scored function $\rho:\Omega\to\{0,1,2,3,4,6\}$ is defined by $\rho(W)=0$ and $\rho(o)=o$ for $o\neq W$; both a wicket and a dot ball therefore contribute zero runs to the batting total. The outcome $o=5$ runs (an all-run five) is omitted from $\Omega$: it occurs in fewer than 0.1% of deliveries in the historical data and its inclusion has negligible effect on computed probabilities. Wides and no-balls are handled at the data-preprocessing stage and are not included in the legal delivery count.
3.2 Batting MDP
For the batting side chasing a target, define the win probability value function:
$V^{B}(r,b,w)=\mathbb{P}\!\left(\text{batting side wins}\;\middle|\;\text{state }(r,b,w),\;\text{order }\sigma\right)$
Terminal conditions: The batting side wins if and only if the target is reached:
$V^{B}(r,b,w)=\begin{cases}1&\text{if }r\leq 0\\[4.0pt] 0&\text{if }b=0\text{ or }w=0,\text{ and }r>0\end{cases}$
Transition function: Given on-strike batsman $i$ in phase $\phi(b)$, outcome $o$ occurs with probability $p_{i}^{o,\phi}$:
$T\!\left((r,b,w),\,o\right)=\begin{cases}(r,\;b-1,\;w-1)&o=W\\[3.0pt] (r-\rho(o),\;b-1,\;w)&o\in\{0,1,2,3,4,6\}\end{cases}$
Bellman equation:
$\boxed{V_{t}^{B}(r,b,w)=\sum_{o\in\Omega}p_{i}^{o,\phi(b)}\cdot V_{t-1}^{B}\!\left(T(r,b,w,o)\right)}$
where $t=b$ denotes the number of balls remaining, so t1t-1 refers to the state after the current delivery is bowled. Expanding over all outcomes:
$V_{t}^{B}(r,b,w)=p_{i}^{W,\phi}\cdot V_{t-1}^{B}(r,\,b-1,\,w-1)+\sum_{\rho\in\{0,1,2,3,4,6\}}p_{i}^{\rho,\phi}\cdot V_{t-1}^{B}(r-\rho,\,b-1,\,w)$
Strike rotation is enforced: the on-strike batsman rotates to the non-striker when $\rho(o)$ is odd, or when $b$ is the last ball of an over. This is consequential: which batsman faces the first ball of the Death overs (16–20) depends on the parity of singles accumulated during overs 14–15, and the SR differential between a Death-over specialist and a lower-order batsman can exceed 50 points.
Optimal batting order:
$\sigma^{*}=\arg\max_{\sigma\in\text{Perm}(\mathcal{P})}V^{B}_{\sigma}(r_{0},b_{0},w_{0})$
where $\mathcal{P}$ is the set of remaining batsmen and $\text{Perm}(\mathcal{P})$ is the set of all $|\mathcal{P}|!$ orderings.
3.3 Bowling MDP
For the bowling side defending a total, define the defend probability value function under plan $\pi$:
$V^{\pi}(d,b,w)=\mathbb{P}\!\left(\text{bowling side defends}\;\middle|\;\text{state }(d,b,w),\;\text{plan }\pi\right)$
Terminal conditions:
[…]
where $w_{\max}$ is the number of wickets in hand for the batting side at the intervention point (not necessarily 10 if wickets have already fallen).
Transition function: Given bowler $j=\pi_{k}$ (assigned to the current over $k$) in phase $\phi(b)$:
$T\!\left((d,b,w),\,o\right)=\begin{cases}(d,\;b-1,\;w+1)&o=W\\[3.0pt] (d-\rho(o),\;b-1,\;w)&o\in\{0,1,2,3,4,6\}\end{cases}$
Here $w$ counts additional wickets taken from the intervention point; the innings ends when $w$ reaches $w_{\max}$.
Bellman equation:
$\boxed{V_{t}^{\pi}(d,b,w)=\sum_{o\in\Omega}p_{j}^{o,\phi(b)}\cdot V_{t-1}^{\pi}\!\left(T(d,b,w,o)\right)}$
Optimal bowling plan:
$\pi^{*}=\arg\max_{\pi\in\mathcal{F}}V^{\pi}(d_{0},b_{0},w_{0})$
where $\mathcal{F}$ is the feasible plan set (Section 6.1).
3.4 Duality
At the same match state $(r_{0}=d_{0},\,b_{0},\,w_{0})$, the batting win probability and bowling defend probability are exact complements under any fixed batting order $\sigma$ and bowling plan $\pi$:
$V^{B}_{\sigma}(r_{0},b_{0},w_{0})+V^{\pi}(d_{0},b_{0},w_{0})=1$
This follows directly from the fact that a T20 match has no draw: the batting side either reaches the target (win) or does not (lose). The duality is exploited in the case studies to provide a complete audit: improvements to one side translate directly to losses for the other.
3.5 Exact Backward Induction and the Case for Simulation
The tree that motivates simulation. At each delivery the match state $(r,b,w)$ branches into 7 possible next states, one per outcome in $\Omega=\{W,0,1,2,3,4,6\}$. Expanding this fully gives a tree with up to $7^{b}$ leaf nodes. At a typical mid-innings intervention with $b=44$ balls remaining, that is $7^{44}\approx 10^{37}$ paths — completely intractable to enumerate.
The Markov property collapses this tree into a table. Because the value of state $(r,b,w)$ depends only on those three numbers and not on the history of how that state was reached, many branches converge to the same node. Reaching $(r=50,b=30,w=5)$ via a boundary and dots is identical in value to reaching it via five singles. The state space therefore contains at most $R\times B\times W\approx 221\times 120\times 10\approx 265{,}200$ unique entries rather than $7^{44}$ paths — a reduction of over thirty orders of magnitude.
Exact backward induction fills this table bottom-up: terminal values are set first, then $V_{1}^{B}$ is computed from the terminals, then $V_{2}^{B}$ from $V_{1}^{B}$, and so on up to $V_{b}^{B}$. Each entry requires exactly 7 multiplications and additions. For the bowling problem with fixed player identities this is feasible; the state space has at most $D\times B\times W\approx 221\times 44\times 10\approx 97{,}000$ states and the Bellman recursion terminates in polynomial time. However, the bowling plan $\pi$ itself must be searched over the feasible set $\mathcal{F}$; running backward induction for each candidate plan is inefficient. Monte Carlo simulation with a fixed plan provides an unbiased estimate of $V^{\pi}(d_{0},b_{0},w_{0})$ directly without filling the full state table.
For the batting problem, the state must additionally track which specific batsmen remain, adding a combinatorial factor of $2^{n}$ subsets to the state space. For $n=6$ remaining batsmen, this multiplies the state count by $64\times$, rendering exact backward induction impractical on commodity hardware. Monte Carlo simulation with a fixed order $\sigma$ avoids this by forward-sampling complete innings trajectories from the initial state, estimating the value function at $(r_{0},b_{0},w_{0})$ without computing all intermediate states.
4. Player Profile Engine
4.1 Data
Ball-by-ball records from 1,161 IPL matches (seasons 2008–2025) are sourced from Cricsheet.org in CSV format via the yorkr R package (Ganesh, 2016). The dataset comprises 273,735 total delivery records; after filtering for legal deliveries (excluding wides and no-balls) approximately 264,800 legal balls remain for profile estimation. All matches involving the target teams (KKR–MI on 29 March 2026 and GT–PBKS on 31 March 2026) are excluded to prevent look-ahead bias.
4.2 Phase Assignment
Each delivery is assigned to one of three innings phases by the over number (0-indexed):
Phase Overs (0-indexed) Characteristic Conditions
Powerplay (PP) 0 – 5 Fielding restrictions; elevated scoring rates
Middle (MI) 6 – 14 Spin-friendly; dot-ball accumulation
Death (DE) 15 – 19 Yorkers; pull shots; explosive hitting
Phase assignment is deterministic given the over number; in-over ball position is not used.
4.3 Outcome Counts and Wicket Attribution
For each player $i$ and phase $\phi$, raw outcome counts are accumulated over all $n_{i}^{\phi}$ legal deliveries:
$c_{i}^{o,\phi}=\bigl|\{t:\text{player }i\text{ involved in delivery }t,\;\text{phase}(t)=\phi,\;\text{outcome}(t)=o\}\bigr|,\quad o\in\Omega$
Two distinct attribution rules apply depending on whether $i$ is a batsman or bowler.
Batsman wicket attribution: A delivery contributes $o=W$ to batsman $i$’s counts if and only if $\texttt{wicketPlayerOut}(t)=i$. This correctly attributes caught-behind, LBW, stumped, bowled, and caught dismissals to the dismissed batsman, and excludes run-outs of the non-striker (who is not the on-strike batsman for that delivery).
Bowler wicket attribution: A wicket is credited to bowler $j$ if and only if the dismissal type is not a run-out:
$\mathbf{1}_{\text{bowler wicket}}(t)=\mathbf{1}\!\left[\texttt{wicketPlayerOut}(t)\neq\texttt{nobody}\right]\cdot\mathbf{1}\!\left[\texttt{dismissalType}(t)\notin\{\text{run out}\}\right]$
Run-outs reflect fielding and batting decisions, not bowling quality, and are excluded.
4.4 Laplace Smoothing
Players who have never produced a specific outcome in a specific phase (for example, a spinner who has never been hit for six in Middle overs) would receive a zero probability estimate, which would prevent the simulation from sampling that outcome and bias win/defend probability estimates downward. Add-one (Laplace) smoothing removes all zero probabilities:
$\tilde{c}_{i}^{o,\phi}=c_{i}^{o,\phi}+\alpha,\qquad\alpha=1$
$\hat{p}_{i}^{o,\phi}=\frac{\tilde{c}_{i}^{o,\phi}}{\displaystyle\sum_{o^{\prime}\in\Omega}\tilde{c}_{i}^{o^{\prime},\phi}}$
With $|\Omega|=7$ and $\alpha=1$, the maximum distortion to any probability is $7/n_{i}^{\phi}$, negligible for players with more than 100 deliveries in a phase but meaningful for very sparse profiles, which are addressed by the shrinkage estimator below.
4.5 James–Stein Shrinkage for Phase-Sparse Players
Many IPL players have limited exposure in specific phase-role combinations: a seam bowler who has rarely bowled Middle overs, a spinner who has never bowled Death overs, or a new franchise addition with fewer than two full IPL seasons. For such players, the Laplace-smoothed individual estimate $\hat{\mathbf{p}}_{i}^{\phi}$ is a high-variance estimate from few observations. We apply a James–Stein shrinkage estimator (James and Stein, 1961), which blends each player’s individual estimate toward the population-average profile for that phase.
Population average: For phase $\phi$, the population-average outcome distribution (across all players in the dataset) is:
$\bar{\mathbf{p}}^{\phi}=\frac{\sum_{i}\tilde{\mathbf{c}}_{i}^{\phi}}{\sum_{i}\sum_{o\in\Omega}\tilde{c}_{i}^{o,\phi}}$
Blended profile:
$\mathbf{p}_{i}^{\phi}=\lambda_{i}^{\phi}\cdot\hat{\mathbf{p}}_{i}^{\phi}+(1-\lambda_{i}^{\phi})\cdot\bar{\mathbf{p}}^{\phi}$
Data-adaptive blend weight:
$\lambda_{i}^{\phi}=\frac{n_{i}^{\phi}}{n_{i}^{\phi}+n_{\min}},\qquad n_{\min}=50$
When $n_{i}^{\phi}=0$ (no historical data in that phase), $\lambda_{i}^{\phi}=0$ and the profile collapses entirely to the population average, a conservative but unbiased prior. As $n_{i}^{\phi}\to\infty$, $\lambda_{i}^{\phi}\to 1$ and the individual estimate dominates. The threshold $n_{\min}=50$ reflects approximately 8 full overs of data, sufficient for phase-specific distributions to be statistically meaningful. This estimator is a practical analogue of the James–Stein phenomenon: the pooled estimate has lower mean squared error than the individual estimate for any player with fewer than $n_{\min}$ phase-specific deliveries.
4.6 Derived Summary Statistics
From the blended profile $\mathbf{p}_{i}^{\phi}$, the following summary statistics are computed for reporting and diagnostic purposes:
$\text{SR}_{i}^{\phi}=100\cdot\sum_{o\in\Omega}\rho(o)\cdot p_{i}^{o,\phi}\qquad\text{(batsman strike rate)}$
$\text{ER}_{j}^{\phi}=6\cdot\sum_{o\in\Omega}\rho(o)\cdot p_{j}^{o,\phi}\qquad\text{(bowler economy rate, runs per over)}$
$p_{i}^{W,\phi}\qquad\text{(dismissal probability per ball — batsman or bowler)}$
$p_{j}^{0,\phi}\qquad\text{(dot-ball probability per ball — proxy for pressure)}$
The strike rate is the expected runs per ball multiplied by 100; the economy rate is the expected runs per ball multiplied by 6, converting to runs per over. Both are derived from the same underlying probability vector and are therefore consistent with the simulation model: a player’s simulated run output will converge to their profile-implied SR or ER as simulation count increases.
5. Monte Carlo Simulation
5.1 Estimating Win and Defend Probabilities
For a fixed batting order $\sigma$ or bowling plan $\pi$, the value function at the initial state $(r_{0},b_{0},w_{0})$ is estimated by forward simulation of complete innings trajectories:
$\hat{V}=\frac{1}{N}\sum_{n=1}^{N}\mathbf{1}\!\left[\text{trajectory}^{(n)}\text{ results in win/defend}\right]$
This is an unbiased estimator of $V(r_{0},b_{0},w_{0})$ with binomial standard error:
$\text{SE}(\hat{V})=\sqrt{\frac{\hat{V}(1-\hat{V})}{N}}\leq\frac{1}{2\sqrt{N}}$
For N=50, [^0] N=50{,}000 simulations, $\text{SE}(\hat{V})\leq 0.22\%$, sufficient to distinguish orderings or plans differing by 1 pp with a signal-to-noise ratio of at least $4.5\sigma$.
5.2 Simulation Algorithm
Each trajectory is initialised from the intervention state and advanced ball by ball until the innings terminates (target reached, all balls bowled, or batting side all out):
Input: order  (or plan ), initial state (r, b, w)
for each simulation n = 1 to N:
    r  r;  b  b;  w_remaining  w
    striker  [current_position];  non_striker  fixed
    over_ball  0
    while b > 0 and w_remaining > 0 and r > 0:
        phase   phase_of(absolute_over(b))
        bowler j  plan [current_over]       (bowling simulation)
[…]
        if o == W:
            w_remaining -= 1
            striker  non_striker
            non_striker  [next_available]
        else:
            r -= (o)
            if (o) mod 2 == 1 or over_ball == 5:
                swap(striker, non_striker)    # strike rotation
        b -= 1;  over_ball = (over_ball + 1) mod 6
    result[n]  (r  0)    (batting) or (r > 0 and b == 0 or w_remaining == 0)  (bowling)
return mean(result)
5.3 Strike Rotation
Strike rotation is a structurally important feature of the model. After each delivery, the on-strike batsman becomes the non-striker if (a) the runs scored are odd ($\rho(o)\in\{1,3\}$), or (b) the delivery is the final ball of the over. At a wicket, the new batsman (next in the order) becomes the on-striker; the surviving batsman takes the non-striker position. This rule correctly models the alternation of the batting pair and ensures that Death-over specialists are exposed to the appropriate deliveries given realistic run-scoring patterns in preceding overs.
5.4 Vectorised Implementation
All $N$ simulations are advanced in parallel using NumPy vectorised operations. At each ball, $N$ outcomes are sampled in a single call to numpy.random.Generator.choice with pre-built probability arrays; state vectors runs_scored, wickets, and done are updated element-wise. Simulations that have terminated (target reached or all out) are masked from subsequent updates. The full N=50, [^0] N=50{,}000-trajectory evaluation completes in approximately 0.15 seconds per configuration on commodity hardware (Apple M-series processor), enabling the broad search described in Section 6.
A fast-evaluation pass uses $N_{\text{fast}}=5{,}000$ simulations for initial screening during the SA search, reducing per-evaluation time to $\approx 15$ ms, with a final high-precision pass at N=30, [^0] N=30{,}000 for the top-10 candidates identified by SA.
6. Search Algorithms
6.1 Batting Order: Exhaustive Enumeration with Two-Pass Refinement
For a batting pool of size $|\mathcal{P}|=n$, the decision at the intervention point is the complete ordering of all $n$ remaining batsmen in positions $3,4,\ldots,n+2$. The full search space has $n!$ elements. For $n\leq 6$, $n!\leq 720$; for $n\leq 4$, $n!\leq 24$, both tractable with Monte Carlo simulation.
A two-pass strategy is employed:
  • Pass 1 (screening): Evaluate all $n!$ orderings with $N_{1}=3{,}000$ simulations. Retain the top-$K=10$ candidates ranked by $\hat{V}$.
  • Pass 2 (refinement): Re-evaluate the top-$K$ candidates with $N_{2}=20{,}000$ simulations. Report $\sigma^{*}$ as the candidate maximising $\hat{V}$ after Pass 2.
The two-pass strategy reduces total simulation count by approximately $5\times$ relative to evaluating all permutations at $N_{2}$ precision, at negligible risk of misranking the true optimum (the probability that a non-top-$K$ permutation has true value exceeding the top-$K$ by more than the Pass 1 SE is less than $10^{-3}$ for typical problem instances).
6.2 Bowling Plan: Feasibility Constraints
The feasible set $\mathcal{F}$ for bowling plans is defined by two constraints derived from the Laws of Cricket:
Quota constraint: Each bowler $j$ may not bowl more than their remaining over quota:
$\sum_{k=1}^{m}\mathbf{1}[\pi_{k}=j]\leq q_{j}\quad\forall j\in\mathcal{B}$
where $m$ is the number of remaining overs and $q_{j}=4-(\text{overs already bowled by }j)$.
Consecutive-over rule: The same bowler cannot bowl consecutive overs; overs are bowled from alternate ends of the pitch and the same bowler cannot therefore bowl two overs in a row:
$\pi_{k}\neq\pi_{k+1}\quad\forall k\in\{1,\ldots,m-1\}$
A bowling plan violating either constraint is illegal and must be excluded from the search. The consecutive-over rule eliminates approximately 15–20% of otherwise quota-valid plans in typical T20 scenarios.
6.3 Bowling Plan: Simulated Annealing
The feasible set $\mathcal{F}$, even after applying quota and consecutive-over constraints, contains several thousand plans for a typical T20 end-game scenario with six bowlers and eight over slots. Exhaustive high-precision evaluation of all such plans would require $|\mathcal{F}|\times N\approx 5{,}000\times 50{,}000=2.5\times 10^{8}$ simulation steps, which would be prohibitively expensive. Simulated annealing (Kirkpatrick et al., 1983) provides an efficient heuristic that finds near-optimal plans with high probability.
Objective: Maximise $\hat{V}^{\pi}(d_{0},b_{0},w_{0})$ over $\pi\in\mathcal{F}$.
Acceptance rule: A candidate neighbour plan $\pi^{\prime}$ is accepted over the current plan $\pi$ according to the Metropolis criterion:
$A(\Delta V,T)=\begin{cases}1&\text{if }\Delta V>0\\[4.0pt] \exp\!\left(\dfrac{\Delta V}{T}\right)&\text{if }\Delta V\leq 0\end{cases}$
where $\Delta V=\hat{V}^{\pi^{\prime}}-\hat{V}^{\pi}$ and $T$ is the current temperature.
Cooling schedule: Linear cooling from $T_{0}=0.05$ to $\epsilon=10^{-6}$:
$T(\text{step})=T_{0}\cdot\!\left(1-\frac{\text{step}}{N_{\text{steps}}}\right)+\epsilon,\qquad N_{\text{steps}}=8{,}000$
Neighbourhood function: A neighbour $\pi^{\prime}$ is generated by selecting a uniformly random over slot $k\in\{1,\ldots,m\}$ and replacing its bowler $\pi_{k}$ with a bowler drawn uniformly from the feasible candidate set:
$\mathcal{C}_{k}\!\left(\pi\right)=\left\{j\in\mathcal{B}\;:\;j\neq\pi_{k},\;\;\text{used}_{\pi}(j)+\mathbf{1}[j=\pi_{k}]-\mathbf{1}[j=\pi_{k}]<q_{j},\;\;j\neq\pi_{k-1},\;\;j\neq\pi_{k+1}\right\}$
This formulation explicitly enforces both quota validity and the consecutive-over rule on every proposed neighbour, ensuring $\pi^{\prime}\in\mathcal{F}$ whenever $\mathcal{C}_{k}(\pi)\neq\emptyset$.
Execution: The SA runs for $N_{\text{steps}}=8{,}000$ iterations, each evaluating a proposed neighbour with $N_{\text{fast}}=5{,}000$ simulations. All unique plans encountered during the search are cached; the top-10 unique plans are then re-evaluated with $N_{\text{refine}}=30{,}000$ simulations for final reporting.
7. Case Study 1: KKR vs MI, 29 March 2026
7.1 Match Context and Intervention Point
The IPL match at Wankhede Stadium saw KKR post a target of 221 runs (220 scored). At ball 2nd.11.6, the final delivery of over 12, MI’s Rohit Sharma was dismissed for 72 runs off 40 balls, caught off the bowling of VG Arora. At the moment of dismissal:
Parameter Value
MI score 148/1
Runs needed ($r_{0}=d_{0}$) 73
Legal balls remaining ($b_{0}$) 44
Required run rate 9.95 per over
MI batsmen at crease RD Rickelton (non-striker, fixed)
MI batting pool $\mathcal{P}$ SA Yadav, Tilak Varma, HH Pandya, Naman Dhir
Both problems, namely the batting order and bowling plan, arise simultaneously from the same intervention state, providing a natural experiment in which the same match state is audited from both perspectives.
7.2 Player Profiles
Table 1: MI Batsman Profiles (Blended, Phase-Specific)
Batsman Hist. balls (MI) Hist. balls (DE) Death SR $p^{W}_{\text{DE}}$ Middle SR $p^{W}_{\text{MI}}$
RD Rickelton 342 156 156.4 0.079 152.9 0.100
SA Yadav 1,765 460 181.3 0.089 143.8 0.035
Tilak Varma 1,132 388 185.5 0.070 137.3 0.030
HH Pandya 844 512 171.6 0.070 134.0 0.037
Naman Dhir 498 204 203.9 0.070 145.9 0.049
The critical structural feature is the phase asymmetry of Naman Dhir: his Death SR (204) exceeds any other candidate by at least 18 points, yet his Middle SR (146) is comparable to SA Yadav’s. This makes him uniquely valuable in the Death overs but risky to promote early. SA Yadav is the optimal bridge through the Middle overs: his dismissal probability (0.035) is the lowest of all candidates, and his Middle-over strike rate (143.8) is the highest among the non-Death specialists. The combination of low wicket risk and competitive scoring rate in the Middle phase makes him the best choice to carry the innings to the Death overs while preserving Tilak Varma and Naman Dhir, both superior Death-over hitters, for the final assault.
Figure 1: Figure 1: MI batsman phase-specific profiles (KKR vs MI, over 12 intervention). The phase asymmetry of Naman Dhir is the key structural feature driving the optimal batting order.
7.3 Batting Order Results
Table 3: Best Achievable Win% by Next-In Choice (Position 3)
Next-In Best Win% Optimal Remaining Order
SA Yadav (actual) 56.5% Naman Dhir → Tilak Varma → HH Pandya
Naman Dhir 55.7% SA Yadav → Tilak Varma → HH Pandya
Tilak Varma 55.1% Naman Dhir → SA Yadav → HH Pandya
HH Pandya 53.3% Naman Dhir → Tilak Varma → SA Yadav
The actual decision to send SA Yadav in immediately at Position 3 was the correct choice. His low Middle-over dismissal probability ($p^{W}_{\text{MI}}=0.035$, lowest of all candidates) and competitive Middle-over strike rate (143.8) make him the optimal bridge through overs 12–15 while preserving Tilak Varma and Naman Dhir for the Death-over assault.
Table 4: All 6 Orderings for Positions 4–6 (SA Yadav at Position 3)
Rank Position 4 Position 5 Position 6 Win%
1 (opt) Naman Dhir Tilak Varma HH Pandya 56.5%
2 Naman Dhir HH Pandya Tilak Varma 56.1%
3 Tilak Varma Naman Dhir HH Pandya 55.2%
4 HH Pandya Naman Dhir Tilak Varma 52.9%
5 (actual) Tilak Varma HH Pandya Naman Dhir 52.4%
6 HH Pandya Tilak Varma Naman Dhir 51.8%
The match order (rank 5) was the second-worst of all six possibilities. Figure 2 shows the best achievable win probability as a function of the next-in choice at Position 3.
Figure 2: Figure 2: Best achievable win probability by next-in choice at Position 3. SA Yadav’s low Middle-over dismissal probability makes him the optimal bridge.
The optimal order promotes Naman Dhir to Position 4, ensuring maximum exposure of his Death SR (204) to the final overs. In the actual match, Dhir entered at Position 6 and faced only 2 balls, both in over 19. The expected additional runs from promoting Dhir two positions can be approximated as:
$\Delta R\approx\Delta b_{\text{Dhir}}\times\frac{\text{SR}_{\text{Dhir}}^{\text{DE}}-\text{SR}_{\text{Varma}}^{\text{DE}}}{100}\approx\Delta b_{\text{Dhir}}\times 0.184$
Even 5 additional Death balls yields approximately 0.92 expected additional runs, sufficient given the proximity to the target, to shift win probability by the observed 4.1 pp.
7.4 Summary
Scenario Win% (MI)
Worst possible batting order 50.0%
Actual batting order 52.4%
Optimal batting order 56.5%
Gain from optimal order +4.1 pp
8. Case Study 2: GT vs PBKS, 31 March 2026
8.1 Match Context and Intervention Point
Gujarat Titans posted 162 in their first innings. At ball 2nd.9.3, the third delivery of over 10 (0-indexed over 9), Punjab Kings’ opener P Simran Singh was caught off Rashid Khan’s bowling. At the moment of dismissal:
Parameter Value
PBKS score 83/2
Runs needed ($r_{0}=d_{0}$) 80
Legal balls remaining ($b_{0}$) 60 (overs 10–19 after Rashid completes over 9)
PBKS wickets in hand 8
Required run rate 8.00 per over
Rashid Khan was mid-over at the moment of dismissal and was committed to completing over 9 (3 remaining balls). The optimisation covers the 10 complete overs that follow (overs 10–19, 0-indexed).
8.2 GT Bowling Resources and Profiles
Table 7: GT Bowler Profiles and Quotas (After Over 9 Completes)
Bowler Overs bowled Quota left Hist. balls (MI) Hist. balls (DE) Middle ER $p^{W}_{\text{MI}}$ Death ER $p^{W}_{\text{DE}}$
Ashok Sharma 1 3 0 0 7.49† 0.043† 9.37† 0.083†
K Rabada 2 2 466 599 7.35 0.053 9.36 0.104
Mohammed Siraj 2 2 451 673 7.35 0.048 9.55 0.067
Rashid Khan 3* 1 2,223 564 6.58 0.047 8.40 0.075
Washington Sundar 2 2 654 81 7.07 0.033 9.61 0.068
M Prasidh Krishna 0 4 398 505 7.47 0.047 9.79 0.086
*Rashid Khan bowled 3 full overs plus 3 balls of over 9; quota = 1 full remaining over.
†Ashok Sharma profile = population average ($\lambda=0$); zero historical IPL deliveries in dataset.
Figure 3: Figure 5: GT bowler landscape at the over 10 intervention point. The optimal Death-over bowler (Rashid Khan) is in the top-left; the most expensive deployed bowler (M Prasidh Krishna) is toward the bottom-right.
Key features of this profile table: - Rashid Khan has the lowest economy in both Middle (6.58) and Death (8.40) phases, making him the most valuable bowler, but his remaining quota is only 1 over. - Washington Sundar shows a sharp phase gradient: Middle ER 7.07 (competitive) vs Death ER 9.61 (second-most expensive). He should not bowl Death overs. - M Prasidh Krishna has the highest Death economy (9.79) of any GT bowler, yet has 4 overs of remaining quota. - Mohammed Siraj was not used in the actual plan despite 2 overs of quota and competitive profiles in both phases (Middle ER 7.35, Death ER 9.55).
Figure 4: Figure 6: Best achievable defend probability by over 10 opener choice. Siraj’s two-over quota makes him the highest-value opener despite not being unused in the actual plan.
8.3 Actual and Optimal Bowling Plans
Table 8: Actual GT Bowling Plan (Overs 10–19)
Over Bowler Phase Economy
10 Ashok Sharma MI 7.49
11 Rashid Khan MI 6.58
12 M Prasidh Krishna MI 7.47
13 Washington Sundar MI 7.07
14 M Prasidh Krishna MI 7.47
15 K Rabada DE 9.36
16 M Prasidh Krishna DE 9.79
17 Ashok Sharma DE 9.37
18 M Prasidh Krishna DE 9.79
19 Washington Sundar DE 9.61
Actual defend% 39.1%
Table 9: Optimal GT Bowling Plan (Overs 10–19)
Over Bowler Phase Economy Changed?
10 Mohammed Siraj MI 7.35
11 Washington Sundar MI 7.07
12 K Rabada MI 7.35
13 Mohammed Siraj MI 7.35
14 Washington Sundar MI 7.07
15 Ashok Sharma DE 9.37
16 Rashid Khan DE 8.40
17 Ashok Sharma DE 9.37
18 K Rabada DE 9.36
19 Ashok Sharma DE 9.37
Optimal defend% 44.3%
Table 10: Top-5 GT Bowling Plans
Rank Ov10 Ov11 Ov12 Ov13 Ov14 Ov15 Ov16 Ov17 Ov18 Ov19 Defend%
1 (opt) Siraj Sundar Rabada Siraj Sundar Ashok Rashid Ashok Rabada Ashok 44.3%
2 Sundar Krishna Sundar Siraj Rabada Ashok Rashid Ashok Rabada Ashok 44.0%
3 Sundar Siraj Sundar Rabada Siraj Ashok Rashid Ashok Rabada Ashok 43.9%
4 Siraj Sundar Siraj Krishna Rashid Ashok Rabada Ashok Rabada Ashok 43.5%
Actual Ashok Rashid Krishna Sundar Krishna Rabada Krishna Ashok Krishna Sundar 39.1%
Figure 5: Figure 7: Top 10 GT bowling plans heatmap. The vertical divider separates Middle overs (left) from Death overs (right); Rashid Khan appears in Death over 16 in all top plans.
Three structural errors characterise the actual plan:
  1. 1.
    Rashid Khan deployed in Middle (over 11): His advantage over the best alternative (Washington Sundar, ER 7.07) in Middle is 0.49 RPO. His advantage over M Prasidh Krishna (ER 9.79) in Death is 1.39 RPO, nearly three times as large. Deploying Rashid in Middle over 11 rather than Death over 16 forfeited this asymmetry.
  2. 2.
    M Prasidh Krishna used for 4 overs, including 2 Death overs (16, 18): Krishna’s Death ER (9.79) is the highest of any GT bowler. His 2 Death overs are expected to concede approximately 1.4 additional runs each relative to the optimal assignment, approximately 2.8 runs cumulatively.
  3. 3.
    Mohammed Siraj not used: With 2 overs of quota and competitive Middle (7.35) and Death (9.55) profiles, Siraj’s omission was the largest single resource-allocation error. The optimal plan uses him in overs 10 and 13 (both Middle phase).
Figure 6: Figure 8: Actual vs optimal GT bowling plan (overs 10–19). The total gain of +5.2 $p$ is distributed across multiple substitutions, with Siraj’s inclusion and Rashid’s repositioning contributing the most.
8.4 Statistical Significance
The $z$-score for the bowling gap in the GT vs PBKS case:
$z=\frac{0.052}{\sqrt{2}\times 0.00187}\approx 19.7$
This is far beyond any conventional significance threshold, confirming that the observed 5.2 pp gap is a stable structural property of the plan, not simulation variance.
9. Comparative Analysis and Discussion
9.1 The Phase-Specificity of Bowling Value
The GT vs PBKS case illustrates the core finding: aggregate economy rate is a poor proxy for in-match bowling value. M Prasidh Krishna has an acceptable overall economy, yet his Death-over economy (9.79 RPO) is the highest of any GT bowler. This divergence is invisible in aggregate statistics but is fully captured by the phase-specific profile. The primary misallocations in the GT plan are summarised in Table 11.
Table 11: Primary Misallocations — GT vs PBKS
Bowler misused Phase deployed Death ER Better option Better Death ER Cost (RPO)
MP Krishna Death (2 overs) 9.79 Rashid Khan 8.40 −1.39
Siraj Not used −0.49 per over
Teams should maintain and consult phase-specific profiles, not aggregate metrics, when making in-match bowling decisions. The difference between using M Prasidh Krishna and Rashid Khan in a Death over is 1.39 expected runs — across two overs that is nearly 3 runs, decisive at a margin of 80 to defend in 10 overs.
9.3 The Rashid Khan Problem: Scarcity of the Best Resource
The GT vs PBKS case illustrates a generalised decision problem absent from the KKR case: optimal allocation of a scarce high-value resource. Rashid Khan was the best available bowler in both Middle and Death phases, but with only 1 over of quota remaining, the batting captain had to choose where to deploy him. The model’s recommendation, namely Death over 16, reflects the principle that scarce resources should be deployed where their marginal contribution is largest, not merely where they are good. Since Rashid’s margin over alternatives is 1.39 RPO in Death versus 0.49 RPO in Middle, the Death deployment is nearly three times more valuable per over. This principle generalises to any team holding a single over of a premium bowler at the death stage of an innings.
9.4 Batting and Bowling Duality in Practice
The KKR vs MI case provides an audit of the batting decision at the same intervention state where the bowling decision could also have been studied. The batting MDP value function $V^{B}$ and the bowling defend probability $V^{\pi}$ are exact complements at any shared state: $V^{B}+V^{\pi}=1$. The actual batting order (rank 5 of 6) gave MI a win probability of 52.4%; the optimal order raises this to 56.5%. The same duality means that a bowling optimisation at that state would directly reduce MI’s win probability by an equivalent margin, illustrating how both sides of the decision can be audited from the same framework.
9.5 Limitations
Static historical profiles. Player profiles are estimated from all historical IPL deliveries with equal weight. Current form, fatigue, pitch conditions, and opponent-specific tendencies are not captured. A bowler who has been expensive in recent matches, or a batsman in exceptional form, may be systematically misrated. Exponential decay weighting over recent innings, or Bayesian updating within the current match, would partially address this.
Opponent-agnostic profiles. Batsman profiles are averaged across all bowlers faced; bowler profiles are averaged across all batsmen faced. Matchup-specific effects, such as a left-arm spinner facing left-handed versus right-handed batsmen or a pace bowler against an aggressive versus defensive batsman, are real but are not modelled here due to the sparsity of matchup-level data across phases. A tensor decomposition of outcomes over (bowler, batsman, phase) would extend the framework to capture these effects.
Pre-committed plans. The bowling plan $\pi$ is assumed to be chosen once at the intervention point and then executed regardless of how the match state evolves. In practice, the bowling captain can and should revise the plan at each over boundary given new information (runs conceded in the previous over, a wicket taken, a batsman promoted). A fully adaptive policy $\pi:(d,b,w)\to j$ would require solving the full MDP over $(d,b,w,\mathcal{Q})$, the augmented state that includes remaining quotas, which is feasible in principle but increases the state space by a factor of $\prod_{j}(q_{j}+1)$.
Ball-by-ball independence. Consecutive ball outcomes are treated as i.i.d. given the current phase and player identities. Momentum effects such as a batsman in rhythm after a boundary or a bowler tightening their line under pressure are not captured. Incorporating a latent pressure or momentum state via a Hidden Markov Model would extend the framework at the cost of additional estimation complexity.
10. Conclusions
This paper has presented a unified Markov Decision Process framework for optimising two of the most consequential in-match decisions in T20 cricket: the batting order at each wicket fall and the bowling plan across the remaining over slots. The framework is grounded in a three-phase player profile engine with James–Stein shrinkage for data-sparse players, evaluated by vectorised Monte Carlo simulation, and searched by exhaustive enumeration (batting order) and simulated annealing (bowling plan).
Applied to two intervention points from the 2026 IPL season, the main findings are:
  1. 1.
    Batting order (KKR vs MI): The optimal batting order improves MI’s win probability by 4.1 pp (52.4% → 56.5%). The actual order was the second-worst of all six feasible permutations. The primary driver is the phase asymmetry of Naman Dhir: his Death SR (204) is superior to any alternative, but he entered at Position 6 and faced only 2 balls. Promoting him to Position 4 ensures exposure of this advantage to the decisive Death-over deliveries.
  2. 2.
    Bowling plan (GT vs PBKS): The optimal bowling plan improves GT’s defend probability by 5.2 pp (39.1% → 44.3%). M Prasidh Krishna (Death ER = 9.79) was used in 4 overs including 2 Death overs, Mohammed Siraj (2-over quota, competitive profile) was not used at all, and Rashid Khan (best Death bowler, ER = 8.40) was deployed in a Middle over where his margin over alternatives was less than one-third of his Death margin.
  3. 3.
    Phase-specificity as the common thread: In both case studies, the source of suboptimality is phase-agnostic deployment: decisions that appear reasonable by aggregate metrics but are exposed as costly when phase-specific profiles are applied. Published aggregate statistics (overall SR, overall economy) suppress the information that matters most for in-match decision support.
  4. 4.
    Methodological: The simulation-based MDP approach is computationally tractable on commodity hardware (under 5 minutes for a full SA run including profile estimation), suggesting feasibility for pitch-side or dugout deployment during a match. With Monte Carlo SE below 0.22% at N=50, [^0] N=50{,}000, the precision is sufficient to rank orderings or plans separated by 1 pp with high confidence.
Phase-specific quantitative decision support, grounded in a principled stochastic optimisation framework, offers a demonstrably large and actionable improvement over the heuristic approach that currently dominates in-match tactical decision-making.
Acknowledgements
Ball-by-ball data sourced from Cricsheet.org (Stevenson, 2023) under Creative Commons Attribution licence. Analysis conducted in Python 3.11 using NumPy 1.26, pandas 2.1, and matplotlib 3.8. Code and figures are available at: https://github.com/tvganesh/T20-MDP-optimisation.
Appendix: Notation Summary
Symbol Definition
$(r,b,w)$ Match state: runs remaining, balls remaining, wickets
$R_{\max}$ Maximum runs in the state space
$B=120$ Total balls in T20 innings
$\mathcal{P}$ Pool of available batsmen
$\sigma$ Batting order — permutation of $\mathcal{P}$
$\pi=(\pi_{1},\ldots,\pi_{m})$ Bowling plan — ordered over assignments
$\sigma^{*},\;\pi^{*}$ Optimal batting order, optimal bowling plan
$\phi\in\{\text{PP, MI, DE}\}$ Innings phase
$\Omega=\{W,0,1,2,3,4,6\}$ Ball outcome set
$\rho(o)$ Runs scored under outcome $o$
$\mathbf{p}_{i}^{\phi}=(p_{i}^{o,\phi})_{o\in\Omega}$ Outcome probability vector: player $i$, phase $\phi$
$T(\mathcal{S},o)$ State transition function
$V_{t}^{B}(r,b,w)$ Batting win probability at time $t=b$ balls remaining
$V_{t}^{\pi}(d,b,w)$ Bowling defend probability at time $t=b$ balls remaining
$\text{SR}_{i}^{\phi}$ Batsman $i$ strike rate in phase $\phi$
$\text{ER}_{j}^{\phi}$ Bowler $j$ economy rate in phase $\phi$
$p_{i}^{W,\phi}$ Dismissal/wicket probability per ball
$p_{j}^{0,\phi}$ Dot-ball probability per ball
$c_{i}^{o,\phi}$ Raw outcome count for player $i$, outcome $o$, phase $\phi$
$\alpha=1$ Laplace smoothing constant
$\bar{\mathbf{p}}^{\phi}$ Population-average profile for phase $\phi$
$\lambda_{i}^{\phi}$ James–Stein blend weight
$n_{i}^{\phi}$ Historical legal deliveries for player $i$ in phase $\phi$
$n_{\min}=50$ Shrinkage threshold
N=50, [^0] N=50{,}000 Monte Carlo simulation count (final evaluation)
$N_{\text{fast}}=5{,}000$ Monte Carlo count during SA search
$\hat{V}$ Monte Carlo estimate of value function
[…]
$q_{j}$ Remaining over quota for bowler $j$
$\mathcal{F}$ Set of feasible bowling plans
$\mathcal{C}_{k}(\pi)$ Feasible candidate set for over slot $k$
$A(\Delta V,T)$ SA acceptance probability
$T_{0}=0.05$ Initial SA temperature
$N_{\text{steps}}=8{,}000$ SA iteration count